How to solve this? A school play sells adult and child tickets. On Friday they sold 55 child tickets and 20 adult tickets for $200. On Saturday they sold 72 child tickets and 36 adult tickets for $306. What is the cost of each individual ticket?
Answer by kggoyani(2)

Let, price of individual child ticket is $x, and price of individual adult ticket is $y.
Now, on Friday we can write an equation is 55x+20y=200....(1).
and on Saturday we can write an equation is 72x+36y=306.....(2).
so, Equation 1: 55x+20y=200
Equation 2: 72x+36y=306
now, we are going to solve both equastions.
you may use http://www.webmath.com/solver2.html to solve equestions.
after solving you will find x=2 and y=4.5
Hence, we can say that the cost of child ticket is $2, and the cost of adult ticket is $4.5.